# On Simultaneous Information and Energy Transmission through Quantum Channels

### v6

@@ VI.2 Properties of Private Capacity-Power Function
From the properties of degradable channels [3], we know a channel $\mathcal{N}$ is degradable if its complimentary channel $\mathcal{N}^{c}$ can be written as
@@ VI.3.1 Private Capacity-Power Function for Amplitude Damping Channel
We take the Hamiltonian to be $\sigma_{3}$ as we had done for calculating the unassisted capacity-power function for depolarizing channels. To calculate the private capacity-power function we take {|$\{\ket{\psi_{i}}\}$, where |$\ket{\psi_{i}}=\cos(\theta_{i}/2)\ket{0}+e^{i\phi_{i}}\sin(\theta_{i}/2)\ket{1}$, on the Bloch sphere where each state is chosen with probability $\{p_{i}\}$, similar to what we did for the unassisted depolarizing channel. We plot results in Figs. 15 and 16. Clearly, the private capacity-power function is concave in $B$.
Figure 15: Variation of private channel capacity power function $P_{1}^{Q}(B)$ for amplitude damping channel for different values of noise parameter $\lambda$.
Figure 16: Projection of private channel capacity power function $P_{1}^{Q}(B)$ for amplitude damping channel on $B$-$\lambda$ plane.
@@ VI.3.2 Private Capacity-Power Function for Depolarizing Channel
To calculate the private capacity-power function of this channel, we consider an ensemble of 3 random pure qubit states as above with corresponding probabilities and maximized over all such ensembles. The results are plotted in Figs. 17 and 18.
Figure 17: Variation of private channel capacity power function $P_{1}^{Q}(B)$ for depolarizing channel with the channel noise parameter $\lambda$.
Figure 18: Projection of private channel capacity power function $P_{1}^{Q}(B)$ for depolarizing channel on $B$-$\lambda$ plane.
@@ VII Discussion and Comments
It is also of interest to extend Blahut-Arimoto techniques [35, 36, 37, 38, 39, 40, 41] for quantum capacity-power functions to efficiently optimize over states. The introduction of random matrix theory into this channel capacity framework can give us insights and help us to build more understanding of channel capacity without going to rigorous optimization techniques.

### v7

@@ VI.2 Properties of Private Capacity-Power Function
From the properties of degradable channels wilde_book , we know a channel $\mathcal{N}$ is degradable if its complimentary channel $\mathcal{N}^{c}$ can be written as
@@ VI.3.1 Private Capacity-Power Function for Amplitude Damping Channel
We take the Hamiltonian to be $\sigma_{3}$ as we had done for calculating the unassisted capacity-power function for depolarizing channels. To calculate the private capacity-power function we take {|$\{\ket{\psi_{i}}\}$, where |$\ket{\psi_{i}}=\cos(\theta_{i}/2)\ket{0}+e^{i\phi_{i}}\sin(\theta_{i}/2)\ket{1}$, on the Bloch sphere where each state is chosen with probability $\{p_{i}\}$, similar to what we did for the unassisted depolarizing channel. We consider an ensemble of 3 random pure qubit states as we took for unassisted case, with corresponding probabilities and maximized over all such probabilities. We plot results in Figs. 15 and 16. Clearly, the private capacity-power function is concave in $B$.
Figure 15: Variation of private channel capacity-power function $P_{1}^{Q}(B)$ for amplitude damping channel for different values of noise parameter $\lambda$.
Figure 16: Projection of private channel capacity-power function $P_{1}^{Q}(B)$ for amplitude damping channel on $B$-$\lambda$ plane.
@@ VI.3.2 Private Capacity-Power Function for Depolarizing Channel
To calculate the private capacity-power function of this channel, we consider an ensemble of 3 random pure qubit states as we took for unassisted case, with corresponding probabilities and maximized over all such probabilities. The results are plotted in Figs. 17 and 18.
Figure 17: Variation of private channel capacity-power function $P_{1}^{Q}(B)$ for depolarizing channel with the channel noise parameter $\lambda$.
Figure 18: Projection of private channel capacity-power function $P_{1}^{Q}(B)$ for depolarizing channel on $B$-$\lambda$ plane.
@@ VII Discussion and Comments
It is also of interest to extend Blahut-Arimoto techniques Blahut ; arimoto ; effifcient_approx_channel_capa_classical ; effifcient_approx_channel_capa_quantum ; bla_ari_type ; quant_blahut_ari ; eff_bound_unknown_channel for quantum capacity-power functions to efficiently optimize over states. The introduction of random matrix theory into this channel capacity framework can give us insights and help us to build more understanding of channel capacity without going to rigorous optimization techniques.